# Frontiers of Generative AI for Network Optimization: Theories, Limits, and Visions

### v1

@@ Abstract
While interest in the application of generative AI (GenAI) in network optimization has surged in recent years, its rapid progress has often overshadowed critical limitations intrinsic to generative models that remain insufficiently examined in existing literature. This survey provides a comprehensive review and critical analysis of GenAI’s role in network optimization. We focus on the two dominant paradigms of generative models including generative diffusion models (GDMs) and large pre-trained models (LPTMs), and organize our discussion around a categorization we introduce, dividing network optimization problems into two primary formulations: one-shot optimization and Markov decision process (MDP). We first trace key developments, including foundational contributions from the broader AI community, and systematically categorize current efforts in the networking field. We also briefly review frontier applications of GDMs and LPTMs in other related networking tasks, providing additional context of this survey. Building on this, we present theoretical generalization bounds for GDMs in both one-shot and MDP settings, offering insights into the fundamental factors affecting model performance. Most importantly, we reflect on the overestimated perception of GenAI’s general capabilities and caution against the all-in-one illusion it may convey. We highlight critical limitations, including difficulties in satisfying hard constraints, limited concept understanding, and the inherently probabilistic nature of outputs. We also propose key directions for future research, such as bridging the gap between generation and optimization. Although these two tasks are increasingly integrated in practical applications, they differ fundamentally in both objectives and underlying mechanisms, necessitating a deeper understanding of their theoretical connections. Ultimately, this survey aims to provide researchers and practitioners with a structured overview and a deeper insight into the strengths, limitations, and potential of GenAI in network optimization.
Index Terms:
Network optimization, generative AI, diffusion model, large language model, 6G.
@@ I Introduction
Network optimization refers to a set of mechanisms, algorithms, and models designed to monitor, manage, and enhance the performance of communication networks (whether wireless or wired) [5]. The targeted performance metrics may vary across scenarios and typically include latency, coverage, throughput, reliability, and energy efficiency. While some network improvements involve high-level modifications to protocols or architectural mechanisms [6, 7], much of the optimization effort in practice focuses on parameter control and decision-making within existing operational frameworks, especially in the context of near-real-time network management (e.g., in the RIC). In this survey, we concentrate on how Generative AI (GenAI) contributes to this type of practical network optimization, where network control tasks are typically formulated as mathematical optimization problems. These formulations involve: (i) optimization variables, representing the parameters to be adjusted; (ii) input parameters, capturing system states or environmental information; (iii) an objective function, describing the performance goal; and (iv) constraints, encoding operational or physical limits. Depending on the dependency structure between decisions over time, these problems can be modeled either as one-shot optimization, where each decision is made independently, or as a Markov decision process (MDP), where decisions are made sequentially and influenced by state transitions.
@@
Current network optimization tasks frequently address scenarios such as computation offloading [8], throughput maximization in reconfigurable intelligent surface (RIS)-assisted networks [9], RAN slicing for vehicular networks [10], energy-efficient cloud services [11], traffic load balancing [12], and edge-assisted satellite–terrestrial systems [13]. These approaches predominantly fall into two categories: data-driven methods [8, 9, 10] and numerical optimization algorithms [11, 12, 13]. However, with the advent of 6G and its anticipated demands including access network slicing for diverse services, large-scale swarm coordination, and intelligent interpretation of emerging service requirements, traditional methods face growing limitations. Data-driven models often struggle with generalization to unseen scenarios, while numerical techniques face scalability challenges as problem complexity increases.
@@
Against this backdrop, the rapid advancement of GenAI offers promising opportunities for addressing complex network optimization problems in a more scalable and generalizable manner. GenAI, in contrast to discriminative AI, focuses on learning the underlying global probability distribution of data, enabling the generation of new samples through distribution-based sampling (e.g., in image synthesis). Discriminative AI, by comparison, typically aims to learn from labeled data to distinguish between input samples, as in classification or regression tasks. Because GenAI models the full data distribution, it inherently possesses discriminative capabilities and often exhibits strong cross-domain generalization even when its outputs are probabilistic. It also inherits the scalability of deep learning for high-dimensional data and has become increasingly adaptable with the rise of large-scale models. Earlier generative models such as variational auto-encoders (VAEs) [14], generative adversarial networks (GANs) [15], and normalizing flow (NF) [16] primarily evolved from density estimation approaches. More recent frameworks like the generative flow network (GFlowNets) [17] have been developed to handle discrete combinatorial optimization tasks in AI for science (AI4S). However, these models still face limitations in architecture and applicability. Despite ongoing use in areas such as channel estimation and noise modeling [18, 19, 20, 21], they have produced few results in complex network optimization scenarios.
@@
Today, the two most prominent classes of GenAI are generative diffusion models (GDMs) and large language models (LLMs). GDMs, originally introduced for image generation with denoising diffusion probabilistic models (DDPMs) [22] as the representative example, generate samples by iteratively denoising from a known noise distribution. LLMs [23], known for their scalability, have shown remarkable capabilities in natural language processing and reasoning. While LLMs have shown promising results in communication and networking applications, their scope is largely limited to language-centric tasks and fails to fully encompass the diverse modalities (such as channel state information, graph structures, and signal traces) common in network optimization. To better capture this broader range of generative modeling efforts, we adopt the term large pre-trained models (LPTMs) [24] in this survey. The main scope difference of LPTMs compared to LLMs is that LPTMs cover a wider range of modes in communication systems. Accordingly, we treat GDMs and LPTMs as two distinct paradigms throughout the rest of this paper. GDMs refer to models trained using forward-noising and reverse-denoising mechanisms, while LPTMs denote large-scale pre-trained (mainly autoregressive) models built on various types of data beyond natural language, including but not limited to multi-modal and domain-specific representations relevant to networking. Both paradigms support diverse neural architectures and are capable of addressing a wide range of optimization tasks across modalities.
@@
{tblr}
colspec=X[0.2,l] X[0.52,l] X[0.2,l] X[0.52,l], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, Abbreviation & Full term Abbreviation Full term
AI Artificial Intelligence IRS Intelligent Reflecting Surface
AI4S AI for Science ISAC Integrated Sensing And Communication
AIGC AI Generated Content LLM Large Language Model
AoI Age of Information LoRA Low-Rank Adaptation
BBO Black-Box Optimization LPTM Large Pre-Trained Model
BLO Bi-Level Optimization MDP Markov Decision Process
CFG Classifier-Free Guidance MEC Multi-access Edge Computing
CSI Channel State Information MLP Multi-Layer Perceptron
DDIM Denoising Diffusion Implicit Model NF Normalizing Flow
DDPG Deep Deterministic Policy Gradient OOD Out-Of-Distribution
DDPM Denoising Diffusion Probabilistic Model OPRO Optimization by PROmpting
DiT Diffusion Transformer O-RAN Open RAN
DRL Deep Reinforcement Learning QoS Quality of Service
GAN Generative Adversarial Network RAN Radio Access Network
GAT Graph Transformer RIC RAN Intelligent Controller
GCN Graph Convolutional Network RIS Reconfigurable Intelligent Surface
@@
IoT Internet of Things VAE Variational Auto-Encoder
IP Internet Protocol VR Virtual Reality
IQL Implicit Q-Learning vRAN Virtualized RAN
TABLE II: Comparison with relevant surveys or magazines in terms of coverage of scopes.
{tblr}
colspec=X[0.18,c] X[0.32,c] X[0.25,c] X[0.2,c] X[0.2,c] X[0.21,c], hlines, vlines, Reference & Field Network Optimization Covered Consideration of GDM Consideration of LPTM Theoretical Bounds Analysis
[25] Telecom Base Model ○ ○ ● ○
[26] Networking ● ○ ○ ○
[27] Telecom Base Model ● ● ● ○
[28] Telecom Base Model ● ○ ● ○
[29] AI ○ ● ● ○
[30] AI ● ○ ● ○
[31] Networking ● ○ ● ○
[32] Networking ● ● ● ○
[33] Agentic Network ○ ● ● ○
[34] Agentic Network ○ ○ ● ○
[35] Agentic Network ● ● ● ○
[36] Telecom Base Model ○ ○ ○ ○
[37] Agentic Network ●○ ○ ○
[38] Agentic Network ● ● ● ○
[39] Telecom Base Model ○ ○ ● ○
[40] Networking ● ● ○ ○
[41] Telecom Base Model ● ● ● ○
[42] Telecom Base Model ● ● ● ○
[43] Agentic Network ● ● ● ○
Ours Networking ● ● ● ●
@@ I-A Motivation
Many existing surveys and journal articles have not focused specifically on the intersection of GenAI and network optimization [25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Some works place greater emphasis on designing system-level frameworks for applying GenAI [36, 37, 38], rather than examining the construction and underlying mechanisms of GenAI itself. Others tend to treat GenAI (particularly LPTMs) as a general-purpose solution across a wide range of networking tasks [35, 36, 34]. While a number of works provide valuable high-level overviews that reflect on limitations and outline future directions [39, 40, 41, 42, 43], they often fall short of offering in-depth analyses at the theoretical level.
Emerging works from both the networking community [44, 45] and the artificial intelligence (AI) community [46, 47, 48, 49] have begun to reflect critically on the fundamental capabilities and limitations of GenAI when applied to complex reasoning and optimization tasks. Notably, LPTMs often struggle with rule learning, symbolic reasoning, and mathematical accuracy [46, 47, 48, 49, 45]. Transformer-based models face well-documented theoretical bottlenecks in compositionality [50, 51, 52, 53], and GDMs tend to perform poorly in tasks that demand high precision and constraint satisfaction due to their probabilistic generative mechanisms [54, 55]. These findings suggest that existing GenAI architectures are inherently limited in addressing many of the advanced demands of network optimization, such as strict adherence to formal rules, interpretability, and robust generalization under domain constraints. Specifically, the availability of background and precisely labeled data for network optimization is far more limited than the abundance of text, video, and image data in the broader AI domain, which may fundamentally constrain the capabilities of GenAI. Moreover, network optimization problems often involve complex domain knowledge across logic, physics, and mathematics related to communication systems, making it challenging for off-the-shelf GenAI models and neural architectures to be directly applicable. Network optimization also imposes stringent requirements on time and space efficiency, necessitating lightweight and distributed model designs that are compatible with communication network infrastructures.
@@
Figure 1: GenAI in network optimization taxonomy map.
@@ I-B Contribution
In this paper, we focus on the role of GenAI in network optimization. We provide a brief classification of GenAI’s work in network optimization in Fig. 1. Starting from a comprehensive survey of existing works on GenAI for network optimization, we provide a theoretical analysis of its strengths and limitations. As shown in Table II, unlike existing relevant surveys and magazine articles, this paper offers a more in-depth and systematic contribution in terms of coverage and theoretical insight. In particular, we theoretically examine the gap between the practical development of GenAI optimizers and the desired characteristics of network optimization. The specific contributions of this paper are as follows.
@@
    Focusing on GDMs as a representative case, we analyze two primary paradigms for their application to network optimization tasks and establish theoretical generalization boundaries for each. This analysis identifies direct factors that influence model performance, thereby offering practical guidance for future implementations. In parallel, we discuss both the strengths and limitations of GenAI in network optimization.
@@
    Inspired by recent advances across multiple disciplines, we reflect critically and argue that both GDMs and LPTMs still face limitations in their application to network optimization, particularly in terms of output uncertainty and complex task reasoning. We also discuss future directions, including the development of fundamental theories, rule-guided generation, and domain-specific benchmarks.
@@ II-A Network Optimization
Network optimization in communication systems is caused by the pursuit of optimal allocation and control of limited physical resources and optimal estimation of the environment [5]. It has high-dimensional, dynamic, and uncertain characteristics, posing challenges such as high computational complexity, strict real-time demands, difficulty in finding a global optimal solution, and limited generalization ability [2]. Common types include signal recovery, resource allocation, and semantic optimization problems. These problems often involve various mathematical properties, such as non-convex optimization, mixed-integer programming [8], bi-level optimization (BLO) [56], and stochastic optimization.
Based on the influence intensity of adjacent output solutions on the current environment parameters, network optimization problems can be divided into two categories according to the mathematical modeling ideas of the problem. First, one-shot optimization problems typically use supervised learning methods in deep learning or specified numerical algorithms. In this paradigm, the problem usually has a clear optimal solution, but consecutive decisions are weakly correlated. In contrast, MDP problems, which involve stochastic decision-making, are more suited to unsupervised deep reinforcement learning (DRL) methods. These methods enable neural networks to learn optimal policies through interaction with the environment to maximize long-term rewards. It is noteworthy that MDP problems do not admit a single deterministic optimal decision due to the probabilistic transitions between states, instead, the objective is to find a policy that maximizes the expected return. Thus, MDP problems emphasize the strong correlation between consecutive decisions.
As shown in Fig. 3 and Table III, both approaches must meet near-real-time inference demands but differ significantly in training and solution characteristics. One-shot optimization is more effective when a clear optimal solution exists, while DRL policies are better for handling sequential decision-making in dynamic environments.
TABLE III: Comparison of the two main roles of GenAI in network optimization.
{tblr}
colspec=X[0.2,l] X[0.52,c] X[0.3,c] X[0.42,c] X[0.43,c] X[0.35,c], cells = c, hlines, vlines, rowsep = 2pt, row1 = 0ex, Role of GenAI & Problem Modeling Form Typical Training Paradigm Solution Characteristics Correlation between two consecutive decisions Resource Allocation Frequency

### v2

@@ Abstract
While interest in the application of generative AI (GenAI) in network optimization has surged in recent years, its rapid progress has often overshadowed critical limitations intrinsic to generative models that remain insufficiently examined in existing literature. This survey provides a comprehensive review and critical analysis of GenAI’s role in network optimization. We focus on the two dominant paradigms of generative models, including generative diffusion models (GDMs) and large pre-trained models (LPTMs), and organize our discussion around a categorization we introduce, dividing network optimization problems into two primary formulations: one-shot optimization and Markov decision process (MDP). We first trace key developments, including foundational contributions from the broader AI community, and systematically categorize current efforts in networking. We also briefly review frontier applications of GDMs and LPTMs in other related networking tasks, providing additional context for this survey. Building on this, we present theoretical generalization bounds for GDMs in both one-shot and MDP settings. The established bounds also enable an equivalent analysis of LPTMs. Most importantly, we reflect on the overestimated perception of GenAI’s general capabilities and caution against the all-in-one illusion it may convey. We highlight critical limitations, including difficulties in satisfying hard constraints, limited concept understanding, and the inherently probabilistic nature of outputs. We also propose key directions for future research, such as bridging the gap between generation and optimization. Although these two tasks are increasingly integrated in practical applications, they differ fundamentally in both objectives and underlying mechanisms, necessitating a deeper understanding of their theoretical connections. Ultimately, this survey aims to provide researchers and practitioners with a structured overview and a deeper insight into the strengths, limitations, and potential of GenAI in network optimization.
@@ I Introduction
Network optimization refers to a set of mechanisms, algorithms, and models designed to monitor, manage, and enhance the performance of communication networks (whether wireless or wired) [5]. The targeted performance metrics may vary across scenarios and typically include latency, coverage, throughput, reliability, and energy efficiency. While some network improvements involve high-level modifications to protocols or architectural mechanisms [6, 7], much of the optimization effort in practice focuses on parameter control and decision-making within existing operational frameworks, especially in the context of near-real-time network management (e.g., in the RIC). In this survey, we concentrate on how Generative AI (GenAI) contributes to this type of practical network optimization, where network control tasks are typically formulated as mathematical optimization problems. These formulations involve: (i) optimization variables, representing the parameters to be adjusted; (ii) input parameters, capturing system states or environmental information; (iii) an objective function, describing the performance goal; and (iv) constraints, encoding operational or physical limits. Depending on the temporal dependency structure of decisions, these problems can be modeled either as one-shot optimization, where each decision is made independently, or as a Markov decision process (MDP), where decisions are made sequentially and influenced by state transitions.
@@
Current network optimization tasks frequently address scenarios such as computation offloading [8], throughput maximization in reconfigurable intelligent surface (RIS)-assisted networks [9], RAN slicing for vehicular networks [10], energy-efficient cloud services [11], traffic load balancing [12], and edge-assisted satellite–terrestrial systems [13]. These approaches predominantly fall into two categories: data-driven methods [8, 9, 10] and numerical optimization algorithms [11, 12, 13]. However, with the advent of 6G and its anticipated demands including access network slicing for diverse services, large-scale swarm coordination, and intelligent interpretation of emerging service requirements, traditional methods face growing limitations. Traditional data-driven models often struggle with generalization to unseen scenarios, while numerical techniques face scalability challenges as problem complexity increases.
@@
Against this backdrop, the rapid advancement of GenAI offers promising opportunities for addressing complex network optimization problems in a more scalable and generalizable manner. GenAI, in contrast to discriminative AI, focuses on learning the underlying global probability distribution of data, enabling the generation of new samples through distribution-based sampling (e.g., in image synthesis). Discriminative AI, by comparison, typically aims to learn from labeled data to distinguish between input samples, as in classification or regression tasks. Because GenAI models the full data distribution, they inherently possess discriminative capabilities and often exhibit strong cross-domain generalization, even when their outputs are probabilistic. It also inherits the scalability of deep learning for high-dimensional data and has become increasingly adaptable as large-scale models have emerged. Earlier generative models, such as variational auto-encoders (VAEs) [14], generative adversarial networks (GANs) [15], and normalizing flow (NF) [16], primarily evolved from density estimation approaches. More recent frameworks, such as the generative flow network (GFlowNets) [17], have been developed to handle discrete combinatorial optimization tasks in AI for science (AI4S). However, these models still face limitations in architecture and applicability. Despite ongoing use in areas such as channel estimation and noise modeling [18, 19, 20, 21], they have produced few results in complex network optimization scenarios.
@@
Today, the two most prominent classes of GenAI are generative diffusion models (GDMs) and large language models (LLMs). GDMs, originally introduced for image generation with denoising diffusion probabilistic models (DDPMs) [22] as the representative example, generate samples by iteratively denoising from a known noise distribution. LLMs [23], known for their scalability, have shown remarkable capabilities in natural language processing and reasoning. While LLMs have shown promising results in communication and networking applications, their scope is largely limited to language-centric tasks and fails to fully encompass the diverse modalities (such as channel state information, graph structures, and signal traces) common in network optimization. To better capture this broader range of generative modeling efforts, we adopt the term LPTMs [24] in this survey. The main scope difference between LPTMs and LLMs is that LPTMs cover a wider range of communication modes. Accordingly, we treat GDMs and LPTMs as two distinct paradigms throughout the rest of this paper. GDMs refer to models trained using forward-noising and reverse-denoising mechanisms, while LPTMs denote large-scale pre-trained (mainly autoregressive) models trained on various types of data beyond natural language, including, but not limited to, multimodal and domain-specific representations relevant to networking. Both paradigms support diverse neural architectures and are capable of addressing a wide range of optimization tasks across modalities. Tables II and III provide a more intuitive comparison of GDMs, LPTMs, and traditional methods, clearly illustrating the distinctive characteristics of GDMs and LPTMs, as well as their respective advantages over traditional approaches.
@@
Abbreviation Full term Abbreviation Full term
AI Artificial Intelligence ISAC Integrated Sensing And Communication
AI4S AI for Science LLM Large Language Model
AIGC AI Generated Content LoRA Low-Rank Adaptation
AoI Age of Information LPTM Large Pre-Trained Model
BBO Black-Box Optimization MDP Markov Decision Process
BLO Bi-Level Optimization MEC Multi-access Edge Computing
CFG Classifier-Free Guidance MLP Multi-Layer Perceptron
CSI Channel State Information NF Normalizing Flow
DDIM Denoising Diffusion Implicit Model OOD Out-Of-Distribution
DDPG Deep Deterministic Policy Gradient OPRO Optimization by PROmpting
DDPM Denoising Diffusion Probabilistic Model O-RAN Open RAN
DiT Diffusion Transformer QoS Quality of Service
DRL Deep Reinforcement Learning RAN Radio Access Network
GAN Generative Adversarial Network RIC RAN Intelligent Controller
GAT Graph Transformer RIS Reconfigurable Intelligent Surface
GCN Graph Convolutional Network RLVR RL with Verifiable Rewards
@@
IoT Internet of Things UE User Equipment
IP Internet Protocol VAE Variational Auto-Encoder
IQL Implicit Q-Learning VR Virtual Reality
IRS Intelligent Reflecting Surface vRAN Virtualized RAN
TABLE II: Comparative summary of GDMs and LPTMs in our taxonomy.
Dimension GDMs LPTMs
Training Objective Denoising score matching Next-token prediction
Generative Mechanism Iterative denoising via stochastic differential processes Direct autoregressive decoding
Data Modality Continuous or discrete signals (solutions, images, trajectories, graphs, etc.) Discrete tokens (language, code, multi-modal pairs)
Model Scale Medium or small-scale (mostly <<10B parameters) Large-scale (>>10B parameters)
Typical Applications Solution synthesis, trajectory planning, channel-signal generation Text-based reasoning, multi-modal understanding and decision-making
TABLE III: Comparison among GDM-based, LPTM-based and traditional (mathematical and conventional learning-based) network optimization.
Dimension GDMs LPTMs Traditional (mathematical and conventional learning-based)
Building Cost Medium High Medium
Inference Latency Medium High Low (low-dim) / High (high-dim)
Interpretability Low Medium High
Constraint Awareness Medium Medium High
Scalability Medium High Low
Generalization Medium High Low
TABLE IV: Comparison with relevant surveys or magazines in terms of coverage of scopes.
Reference Field Network Optimization Covered Consideration of GDM Consideration of LPTM Theoretical Bounds Analysis
[25] Telecom Base Model
[26] Networking
[27] Telecom Base Model
[28] Telecom Base Model
[29] AI
[30] AI
[31] Networking
[32] Networking
[33] Agentic Network
[34] Agentic Network
[35] Agentic Network
[36] Telecom Base Model
[37] Agentic Network
[38] Agentic Network
[39] Telecom Base Model
[40] Networking
[41] Telecom Base Model
[42] Telecom Base Model
[43] Agentic Network
Ours Networking
@@ I-A Motivation
Although many existing surveys and journal articles have discussed the intersection of GenAI and network optimization [25, 26, 27, 28, 29, 30, 31, 32, 33, 34], this specific topic has not been their primary focus of research. Some works place greater emphasis on designing system-level frameworks for applying GenAI [36, 37, 38], rather than examining the construction and underlying mechanisms of GenAI itself. Others tend to treat GenAI (particularly LPTMs) as a general-purpose solution across a wide range of networking tasks [35, 36, 34]. While a number of works provide valuable high-level overviews that reflect on limitations and outline future directions [39, 40, 41, 42, 43], they often fall short of offering in-depth analyses at the theoretical level.
Emerging work from both the networking community [44, 45] and the artificial intelligence (AI) community [46, 47, 48, 49] has begun to critically examine the fundamental capabilities and limitations of GenAI when applied to complex reasoning and optimization tasks. Notably, LPTMs often struggle with rule learning, symbolic reasoning, and mathematical accuracy [46, 47, 48, 49, 45]. Transformer-based models face well-documented theoretical bottlenecks in compositionality [50, 51], and GDMs tend to perform poorly in tasks that demand high precision and constraint satisfaction due to their probabilistic generative mechanisms [52, 53]. These findings suggest that existing GenAI architectures are inherently limited in addressing many of the advanced demands of network optimization, such as strict adherence to formal rules, interpretability, and robust generalization under domain constraints. Specifically, the availability of background and precisely labeled data for network optimization is far more limited than the abundance of text, video, and image data in the broader AI domain, which may fundamentally constrain the capabilities of GenAI. Moreover, network optimization problems often involve complex domain knowledge across logic, physics, and mathematics related to communication systems, making it challenging for off-the-shelf GenAI models and neural architectures to be directly applicable. Network optimization also imposes stringent requirements on time and space efficiency, necessitating lightweight and distributed model designs that are compatible with communication network infrastructures.
@@
Figure 1: GenAI in network optimization : taxonomy map.
@@ I-B Contribution
In this paper, we focus on the role of GenAI in network optimization. We provide a brief classification of GenAI’s work in network optimization in Fig. 1. Starting from a comprehensive survey of existing works on GenAI for network optimization, we provide a theoretical analysis of its strengths and limitations. As shown in Table IV, unlike existing relevant surveys and magazine articles, this paper offers a more in-depth and systematic contribution in terms of coverage and theoretical insight. In particular, we theoretically examine the gap between the practical development of GenAI optimizers and the desired characteristics of network optimization. The specific contributions of this paper are as follows.
@@
    Focusing on GDMs as a representative case,we analyze two primary paradigms (One-shot optimization and MDP) for their application to network optimization tasks and establish theoretical generalization boundaries for each. This analysis identifies direct factors that influence model performance, thereby offering practical guidance for future implementations. In parallel, we discuss both the strengths and limitations of GenAI in network optimization.
@@
    Inspired by recent advances across multiple disciplines, we reflect critically and argue that both GDMs and LPTMs still face limitations in their application to network optimization, particularly in terms of output uncertainty and complex task reasoning. With the empirical studies summarized, we also discuss future directions, including the development of fundamental theories, rule-guided generation, and domain-specific benchmarks.
@@ II-A Network Optimization
Network optimization in communication systems is caused by the pursuit of optimal allocation and control of limited physical resources and optimal estimation of the environment [5]. It has high-dimensional, dynamic, and uncertain characteristics, posing challenges such as high computational complexity, strict real-time demands, difficulty in finding a global optimal solution, and limited generalization ability [2]. Common types include signal recovery, resource allocation, and semantic optimization problems. These problems often involve various mathematical properties, such as non-convex optimization [54], mixed-integer programming [8], bi-level optimization (BLO) [55], and stochastic optimization [56].
Based on the influence intensity of adjacent output solutions on the current environment parameters, network optimization problems can be divided into two categories according to the problem’s mathematical modeling approach. First, one-shot optimization problems typically use supervised learning methods in deep learning or specified numerical algorithms. In this paradigm, the problem usually has a clear optimal solution, but consecutive decisions are weakly correlated. In contrast, MDP problems, which involve stochastic decision-making, are more suited to unsupervised deep reinforcement learning (DRL) methods. These methods enable neural networks to learn optimal policies through interaction with the environment, thereby maximizing long-term rewards. It is noteworthy that MDP problems do not admit a single deterministic optimal policy due to probabilistic state transitions. Instead, the objective is to find a policy that maximizes the expected return. Thus, MDP problems emphasize the strong correlation between consecutive decisions.
As shown in Fig. 3 and Table V, both approaches must meet near-real-time inference demands but differ significantly in training and solution characteristics. One-shot optimization is more effective when a clear optimal solution exists, whereas DRL policies are better suited for sequential decision-making in dynamic environments.
TABLE V: Comparison of the two main roles of GenAI in network optimization.
Role of GenAI Problem Modeling Form Typical Training Paradigm Solution Characteristics Correlation between two consecutive decisions Inference Frequency